# Analytic Subspace Routing: How Recursive Least Squares Works in Continual Learning of Large Language Model

### v1

@@ Abstract
Fine-tuning large language models (LLMs) on data of various domains encompassing advanced language-related tasks. Such a process is essentially a continual learning procedure since it is impractical to re-access the pre-trained data in most parties. This naturally attracts catastrophic forgetting of LLMs’ previously learned knowledge, such as general skills. Existing techniques either leverage previous data to replay, leading to extra computational costs, or utilize a single parameter-efficient module to learn the downstream task, constraining new knowledge absorption with interference among different tasks. To handle these issues, we propose an Analytic Subspace Routing (ASR) to address these challenges. For each task, we isolate the learning within a subspace of deep layers’ features via low-rank adaptation, eliminating knowledge interference between different tasks. Additionally, we propose an analytic routing mechanism to properly utilize knowledge learned in different subspaces. Our approach employs Recursive Least Squares to train a multi-task router model, allowing the router to dynamically adapt to incoming data without requiring access to historical data. Subsequently, the router effectively assigns the current task to an appropriate subspace and has a non-forgetting property of previously learned tasks with a solid theoretical guarantee. Experimental results demonstrate that our method achieves near-perfect retention of prior knowledge while seamlessly integrating new information, effectively overcoming the core limitations of existing methods. Our code will be released shortly.
@@ 1 Introduction
In this paper, we propose an Analytic Subspace Routing (ASR) framework for continual learning of LLMs. Unlike common LLMs CL techniques, the proposed ASR creatively adopts a classic recursive least squares (RLS) closed-form solution as an important mean to battle catastrophic forgetting. This comes in a hybrid structure, i.e., with an analytic router with a closed-form solution branching out in an earlier frozen segment of LLM layers, and the remaining layers tuned by LoRA using normal gradient-based technique. For each task, we isolate the learning within a subspace of deep layers’ features via low-rank adaptation, eliminating knowledge interference among different tasks. We trained a specialized LoRA attachment, and use the analytic router to tell which one to load when inference. The key to avoid forgetting is that the analytic router bears the equivalence between continual training one task at a time, and the joint training by putting all-the-task data simultaneously. The main contributions of this work are listed as follow:
  • We introduce the ASR, a novel closed-form subspace routing approach for continual learning in LLMs, facilitating dynamic and adaptable learning of new tasks.
@@
  • We validate the ASR through comprehensive experiments, achieving state-of-the-art performance on the Trace metric [35].
@@ 3.1 Problem Definition
Figure 2: An overview of our proposed method ASR. All the original parameters of the LLM are frozen. (a) The input data enters the first $L_{f}$ layers of the Transformer blocks in the model, and the feature $h_{\leq L_{f}}(\bm{x})$ are obtained. Features will be expanded to a larger dimension, then an analytic router will be trained in the manner of ridge regression. The trained router will perform recursive updates based on AutoCor $\bm{R}_{k}$ (or $\bm{R}_{k}$ and $\bm{Q}_{k}$) and the new coming dataset.(b) The inference module. The trained router accept the feature as input, and is responsible for selecting which LoRA task model to use during the inference process.
@@ 3.5 Recursive Incremental Update
The proof can be found in the Supplementary material LABEL:app:proof.
@@ 4.2 Baselines and Evaluation Metrics
We compare our proposed method, ASR, with ten baseline approaches commonly used in the context of continuous learning:
@@ 4.4.1 Performance on Continual Learning Tasks
To validate the performance of continual learning, we compared the proposed ASR with baseline methods on the TRACE benchmark using two distinct learning orders. Table 2 illustrates that ASR demonstrates competitive performance in both OP and BWT with significant margins when compared to exemplar-based methods. For instance, when assessed with Order1, ASR achieves an OP of 55.69%, surpassing the second exemplar-free method, SepFT, by 8.06%. This trend persists when evaluating with Order2. When comparing with replay-based methods, although replay-based methods take advantage of using exemplars and achieve general good results when compared with exemplar-free methods except ASR, ASR can still outperform the state-of-the-art repaly-based methods. That is, ASR achieves superior result compared with the state-of-the-art SEEKER with both Order1 and Order2 (55.69% and 55.69% v.s. 54.99% and 54.69%). This leading pattern demonstrates the competitiveness of ASR in adapting LLMs to downstream tasks. Furthermore, ASR stands out as the sole method achieving 0 BWT. This observation showcases ASR’s ability to prevent interference between tasks emphasizing its superior knowledge retention. Notably, ASR consistently produces identical results when learning with different orders, while other compared methods are susceptible the order of learning tasks. This observation demonstrates ASR’s robustness and its invariance to task order.
@@ 4.4.2 Ablation study
Ablation on Components in ASR. To provide a systematical analysis on each component in ASR, we compare our method with two other strategies to demonstrate the effect of the isolated subspace training and analytic router. The alternative methodsinvolve employing a single LoRA adapter for multi-task continual learning and utilizing a distinct LoRA adapter for each task while training the router using BP. For LoRA training, we set the rank as 8 while the learning rate as 1e-4; for BP-based router training, we utilize a two-layer MLP with the same expansion size $E$ as AL-based router. As shown in Figure 3, our method demonstrates performance advantage over directly using a single LoRA and training a router with BP. For the methods of single LoRA, at each step, the new task can obtain competitive results while previous tasks encounter performance drop. This observation indicates the knowledge interference across different tasks. For the BP-based router, results display a more erratic pattern. To delve deeper into this phenomenon, we conducted additional experiments on these two routers.
@@
Analysis on Hyperparameters in ASR. As shown in 3, we present the performance of ASR with different feature layer $L_{f}$ and scale-up dimension $E$. Regarding $L_{f}$, lower feature layers (e.g., $L_{f}=2$) exhibit non-zero BWT and lower OP, possibly due to the general nature of these lower features that compromising the router’s accuracy. Performance initially improves with an increase in $L_{f}$, but starts to degrade beyond $L_{f}>4$. This decline can be attributed to the decreased effectiveness of the remaining layers in task-specific learning with deeper feature layers, leading to a decrease in OP.Regarding $E$, it is noticeable that performance increases as $E$ scales up from 5k to 10k when $L_{f}=4$. However, further increases in $E$ do not result in performance enhancements, indicating that choosing $E=10k$ could be sufficient. The configuration of $L_{f}=4,E=10k$ dimensions yields optimal results.
@@ 4.4.3 General Ability after Continual Learning
To validate ASR’s performance of preserving general knowledge after continual learning in downstream tasks, we report the performance of models after CL on general skills benchmarks. It is crucial to emphasize that our ASR utilizes only 1% of the merged validation subset data across all evaluated tasks exclusively for training the router module. This limited dataset is strictly employed to optimize the router’s decision boundaries through constrained gradient updates, directing the general questions to the origin model with parameters entirely isolated from any exposure to these samples.
As shown in Table 4, ASR demonstrates relatively stable performance across various tasks. When compared with exemplar-free method SeqFT, ASR can maintain better general skill. Although the replay-based methods (i.e., Replay and SEEKER) outperforms ASR, the gaps are relative small. This observation demonstrates the effectiveness of ASR to retain general knowledge while adhere to exemplar-free constrain. This phenomenon stems from two methodological characteristics:
@@ 5 Conclusion
In this work, we propose ASR, a novel framework that enables efficient, memory-free, and scalable continual learning for LLMs. By leveraging Recursive Least Squares (RLS) for a dynamic multi-task routing mechanism, ASR eliminates the need for storing or replaying historical data, addressing core limitations of traditional methods. Through a combination of parameter-efficient adaptation using LoRA and an analytic routing approach, ASR achieves near-perfect retention of prior knowledge while seamlessly integrating new tasks, demonstrating robust and efficient performance. Our approach not only overcomes the core limitations of existing methods such as catastrophic forgetting and computational inefficiency, but also provides a solid theoretical foundation for the non-forgetting property, ensuring stability and reliability in continual learning scenarios.

### v2

@@ Abstract
Fine-tuning large language models (LLMs) on data of various domains encompassing advanced language-related tasks. Such a process is essentially a continual learning procedure since it is impractical to re-access the pre-trained data in most parties. This naturally attracts catastrophic forgetting of LLMs’ previously learned knowledge, such as general skills. Existing techniques either leverage previous data to replay, leading to extra computational costs, or utilize a single parameter-efficient module to learn the downstream task, constraining new knowledge absorption with interference among different tasks. To handle these issues, we propose an Analytic Subspace Routing (Any-SSR) to address these challenges. For each task, we isolate the learning within a subspace of deep layers’ features via low-rank adaptation, eliminating knowledge interference between different tasks. Additionally, we propose an analytic routing mechanism to properly utilize knowledge learned in different subspaces. Our approach employs Recursive Least Squares to train a multi-task router model, allowing the router to dynamically adapt to incoming data without requiring access to historical data. Subsequently, the router effectively assigns the current task to an appropriate subspace and has a non-forgetting property of previously learned tasks with a solid theoretical guarantee. Experimental results demonstrate that our method achieves near-perfect retention of prior knowledge while seamlessly integrating new information, effectively overcoming the core limitations of existing methods. Our code is available at https://github.com/ZHUANGHP/Any-SSR
@@ 1 Introduction
In this paper, we propose an Analytic Subspace Routing (Any-SSR) framework for continual learning of LLMs. Unlike common LLMs CL techniques, the proposed Any-SSR creatively adopts a classic recursive least squares (RLS) closed-form solution as an important mean to battle catastrophic forgetting. This comes in a hybrid structure, i.e., with an analytic router with a closed-form solution branching out in an earlier frozen segment of LLM layers, and the remaining layers tuned by LoRA using normal gradient-based technique. For each task, we isolate the learning within a subspace of deep layers’ features via low-rank adaptation, eliminating knowledge interference among different tasks. We trained a specialized LoRA attachment, and use the analytic router to tell which one to load when inference. The key to avoid forgetting is that the analytic router bears the equivalence between continual training one task at a time, and the joint training by putting all-the-task data simultaneously. The main contributions of this work are listed as follow:
  • We introduce the Any-SSR, a novel closed-form subspace routing approach for continual learning in LLMs, facilitating dynamic and adaptable learning of new tasks.
@@
  • We validate the Any-SSR through comprehensive experiments, achieving state-of-the-art performance on the Trace metric [35].
@@ 3.1 Problem Definition
Figure 2: An overview of our proposed method Any-SSR. All the original parameters of the LLM are frozen. (a) The input data enters the first $L_{f}$ layers of the Transformer blocks in the model, and the feature $h_{\leq L_{f}}(\bm{x})$ are obtained. Features will be expanded to a larger dimension, then an analytic router will be trained in the manner of ridge regression. The trained router will perform recursive updates based on AutoCor $\bm{R}_{k}$ (or $\bm{R}_{k}$ and $\bm{Q}_{k}$) and the new coming dataset.(b) The inference module. The trained router accept the feature as input, and is responsible for selecting which LoRA task model to use during the inference process.
@@ 3.5 Recursive Incremental Update
The proof can be found in the Supplementary material A.
@@ 4.2 Baselines and Evaluation Metrics
We compare our proposed method, Any-SSR, with ten baseline approaches commonly used in the context of continuous learning:
@@ 4.4.1 Performance on Continual Learning Tasks
To validate the performance of continual learning, we compared the proposed Any-SSR with baseline methods on the TRACE benchmark using two distinct learning orders. Table 2 illustrates that Any-SSR demonstrates competitive performance in both OP and BWT with significant margins when compared to exemplar-based methods. For instance, when assessed with Order1, Any-SSR achieves an OP of 55.69%, surpassing the second exemplar-free method, SepFT, by 8.06%. This trend persists when evaluating with Order2. When comparing with replay-based methods, although replay-based methods take advantage of using exemplars and achieve general good results when compared with exemplar-free methods except Any-SSR, Any-SSR can still outperform the state-of-the-art repaly-based methods. That is, Any-SSR achieves superior result compared with the state-of-the-art SEEKER with both Order1 and Order2 (55.69% and 55.69% v.s. 54.99% and 54.69%). This leading pattern demonstrates the competitiveness of Any-SSR in adapting LLMs to downstream tasks. Furthermore, Any-SSR stands out as the sole method achieving 0 BWT. This observation showcases Any-SSR’s ability to prevent interference between tasks emphasizing its superior knowledge retention. Notably, Any-SSR consistently produces identical results when learning with different orders, while other compared methods are susceptible the order of learning tasks. This observation demonstrates Any-SSR’s robustness and its invariance to task order.
@@ 4.4.2 Ablation study
Ablation on Components in Any-SSR. To provide a systematical analysis on each component in Any-SSR, we compare our method with two other strategies to demonstrate the effect of the isolated subspace training and analytic router. The alternative methodsinvolve employing a single LoRA adapter for multi-task continual learning and utilizing a distinct LoRA adapter for each task while training the router using BP. For LoRA training, we set the rank as 8 while the learning rate as 1e-4; for BP-based router training, we utilize a two-layer MLP with the same expansion size $E$ as AL-based router. As shown in Figure 3, our method demonstrates performance advantage over directly using a single LoRA and training a router with BP. For the methods of single LoRA, at each step, the new task can obtain competitive results while previous tasks encounter performance drop. This observation indicates the knowledge interference across different tasks. For the BP-based router, results display a more erratic pattern. To delve deeper into this phenomenon, we conducted additional experiments on these two routers.
@@
Analysis on Hyperparameters in Any-SSR. As shown in 3, we present the performance of Any-SSR with different feature layer $L_{f}$ and scale-up dimension $E$. Regarding $L_{f}$, lower feature layers (e.g., $L_{f}=2$) exhibit non-zero BWT and lower OP, possibly due to the general nature of these lower features that compromising the router’s accuracy. Performance initially improves with an increase in $L_{f}$, but starts to degrade beyond $L_{f}>4$. This decline can be attributed to the decreased effectiveness of the remaining layers in task-specific learning with deeper feature layers, leading to a decrease in OP.Regarding $E$, it is noticeable that performance increases as $E$ scales up from 5k to 10k when $L_{f}=4$. However, further increases in $E$ do not result in performance enhancements, indicating that choosing $E=10k$ could be sufficient. The configuration of $L_{f}=4,E=10k$ dimensions yields optimal results.
@@ 4.4.3 General Ability after Continual Learning
To validate Any-SSR’s performance of preserving general knowledge after continual learning in downstream tasks, we report the performance of models after CL on general skills benchmarks. It is crucial to emphasize that our Any-SSR utilizes only 1% of the merged validation subset data across all evaluated tasks exclusively for training the router module. This limited dataset is strictly employed to optimize the router’s decision boundaries through constrained gradient updates, directing the general questions to the origin model with parameters entirely isolated from any exposure to these samples.
As shown in Table 4, Any-SSR demonstrates relatively stable performance across various tasks. When compared with exemplar-free method SeqFT, Any-SSR can maintain better general skill. Although the replay-based methods (i.e., Replay and SEEKER) outperforms Any-SSR, the gaps are relative small. This observation demonstrates the effectiveness of Any-SSR to retain general knowledge while adhere to exemplar-free constrain. This phenomenon stems from two methodological characteristics:
@@ 5 Conclusion
In this work, we propose Any-SSR, a novel framework that enables efficient, memory-free, and scalable continual learning for LLMs. By leveraging Recursive Least Squares (RLS) for a dynamic multi-task routing mechanism, Any-SSR eliminates the need for storing or replaying historical data, addressing core limitations of traditional methods. Through a combination of parameter-efficient adaptation using LoRA and an analytic routing approach, Any-SSR achieves near-perfect retention of prior knowledge while seamlessly integrating new tasks, demonstrating robust and efficient performance. Our approach not only overcomes the core limitations of existing methods such as catastrophic forgetting and computational inefficiency, but also provides a solid theoretical foundation for the non-forgetting property, ensuring stability and reliability in continual learning scenarios.
Supplementary Material
Appendix A Proof of Eqa. (17)
Proof. According to 15:
$\hat{\bm{W}}_{k+1}^{\text{r}}=\bm{R}_{k+1}\bm{Q}_{k+1}=\bm{R}_{k+1}\bm{Q}_{k}+\bm{R}_{k+1}\tilde{h}_{k+1}^{\mathrm{T}}\bm{Y}_{k+1}.$ (26)
Noticed that $\bm{R}_{k+1}\bm{Q}_{k+1}$ can be written as:
$\bm{R}_{k+1}\bm{Q}_{k}$ $=\bm{R}_{k}\bm{Q}_{k}-\bm{R}_{k}\tilde{h}_{k+1}^{\top}\bm{K}\tilde{h}_{k+1}\bm{R}_{k}\bm{Q}_{k}$
$=(\bm{I}-\bm{R}_{k}\tilde{h}_{k+1}^{\top}\bm{K}_{k+1}\tilde{h}_{k+1})\hat{\bm{W}}_{k},$
Where $\bm{K}=(\bm{I}+\tilde{h}_{k+1}\bm{R}_{k}\tilde{h}_{k}^{T})^{-1}$ and $\bm{K}\in\mathbb{R}^{n\times n}$.
Since
$\bm{K}\bm{K}^{-1}=\bm{K}(\bm{I}+\tilde{h}_{k+1}\bm{R}_{k}\tilde{h}_{k+1}^{T})=\bm{I},$ (27)
we have
$\bm{K}=\bm{I}-\bm{K}\tilde{h}_{k+1}\bm{R}_{k}\tilde{h}_{k+1}^{\top}.$ (28)
Therefore
$\bm{R}_{k+1}\bm{Q}_{k}=(\bm{I}-\bm{R}_{k+1}\tilde{h}_{k+1}^{T}\tilde{h}_{k+1})\hat{\bm{W}}_{k}^{\text{R}}.$ (29)
Instantiate 29 into 15, we have
$\hat{\bm{W}}_{k+1}^{\text{r}}=(\bm{I}-\bm{R}_{k+1}\tilde{h}_{k+1}^{T}\tilde{h}_{k+1})\hat{\bm{W}}_{k}^{\text{R}}+\bm{R}_{k+1}\tilde{h}_{k+1}^{\mathrm{T}}\bm{Y}_{k+1}.$ (30)
which completes the proof.